# MRAMG-Bench: A BeyondText Benchmark for Multimodal Retrieval-Augmented Multimodal Generation

### v1

@@ 7. Conclusion
To adapt the growing importance of generating multimodal answers, MRAMG has emerged as a critical task that aligns with real-world user demands. To address the lack of evaluation resources for this task, we present MRAMG-Bench, a meticulously curated dataset containing 4,800 questions across diverse domains, varying difficulty levels, and rich image contexts. We further propose a comprehensive evaluation strategy, incorporating statistical and LLM-based metrics to rigorously assess both retrieval and generation performance. In addition, we introduce a general multimodal generation framework to enable models to produce interleaved text-image responses. Our evaluation of eleven generative models highlights notable limitations in handling challenging datasets and selecting the correct image order, emphasizing the necessity for deeper exploration of the MRAMG task.

### v2

@@ 7. Conclusion
Driven by the growing demand for multimodal answers, MRAMG has emerged as a critical task that aligns with real-world user needs. To address the lack of evaluation resources for this task, we present MRAMG-Bench, a curated benchmark containing 4,800 QA pairs across diverse domains and varying difficulty levels. We further propose a comprehensive evaluation strategy that incorporates statistical and LLM-based metrics to rigorously assess both retrieval and generation performance. Additionally, we introduce a general MRAMG framework to enable models to produce interleaved text-image responses. Our evaluation of 11 popular generative models highlights notable limitations in handling challenging datasets and selecting the correct image order, emphasizing the necessity for deeper exploration of the MRAMG task.
Acknowledgements.
This work is supported by National Natural Science Foundation of China (92470121, 62402016), National Key R&D Program of China (2024YFA1014003), and High-performance Computing Platform of Peking University. Binghui Li is supported by the Elite Ph.D. Program in Applied Mathematics for PhD Candidates in Peking University.